OpenStudy (tropical):

3^-9 ÷ 3^-12 = ? Help?

OpenStudy (kittiwitti1):

\[3^{-9}+3^{-12}=\frac{1}{3^9}+\frac{1}{3^{12}}\]

OpenStudy (tropical):

Thank you

OpenStudy (sshayer):

\[3^{-9}\div 3^{-12}=\frac{ 3^{-9} }{ 3^{-12} }=3^{-9}*3^{12}=3^{-9+12}=3^3=27\]